Andrea divides up her free time in the ratio: games console : reading = 4 : 1.
She has 200 minutes of free time each day. How much time does she spend on each
activity.
Q.1 Answers
Time she spends on games console Time she spends on reading
1
Expert's answer
2020-06-11T20:06:04-0400

Let x be a time she spends on reading.

So, 4x is a time she spends on games console.

We have: "x+4x=200."

"x=\frac{200}{4+1}=40."

"4x=4*40=160".

Time she spends on reading: 40 minutes.

Time she spends on games console: 160 minutes.